Find the curl F given that F x,y,z = 3x2i+2zj-xk


1
Expert's answer
2022-04-15T14:33:22-0400

"P=3x^2,\;R=2z,\;Q=-x."

"curlF=(R_y-Q_z)i+(P_z-R_x)j+(Q_x-P_y)k="

"=(0-0)i+(0-2)j+(-1-0)k=-2j-k."